# EmoEdit: Evoking Emotions through Image Manipulation

### v1

@@ 4.1 Dataset and Evaluation
There are no publicly available datasets designed specifically for the AIM task. We collected 416 real images as a test set for EmoEdit, sourced from user and artist uploads on the Internet [^1] Corresponding author. To ensure greater diversity in AIM, our data comprises positive, negative and neutral images, categorized by a pre-trained emotion classifier on EmoSet [39]. Our method is adaptable to any user-provided images without requiring additional training.
@@
Table 1: Comparisons with the state-of-the-art methods on global editing, local editing and style-based AIM methods. Six metrics are used, encompassing pixel-level, semantic-level, and emotion-level evaluations.
SDEdit [18] 15.41 0.415 0.460 0.639 38.07% 0.220
PnP [30] 14.39 0.435 0.381 0.850 23.83% 0.095
InsDiff [9] 10.74 0.319 0.506 0.796 19.10% 0.059
P2P-zero [22] 13.75 0.451 0.545 0.685 20.24% 0.066
ControlNet [42] 12.04 0.296 0.600 0.687 36.33% 0.215
BlipDiff [15] 8.99 0.249 0.654 0.811 18.07% 0.045
CLVA [8] 12.62 0.398 0.479 0.758 14.22% 0.018
AIF [34] 14.08 0.539 0.492 0.828 12.74% 0.003
EmoEdit 16.01 0.561 0.300 0.832 64.16% 0.363
@@
Given the multiple objectives of AIM, we assess our method based on three aspects: pixel-level (PSNR, SSIM), semantic-level (LPIPS, CLIP-I) and emotion-level (Emo-A, Emo-S). PSNR measures the reconstruction quality by comparing the pixel-by-pixel similarity between the original and edited images, indicating the level of noise and distortion introduced during the manipulation process. SSIM [33] evaluates the perceptual quality of images by comparing low-level structural information, luminance, and contrast. LPIPS [44] leverages deep learning models to assess the perceptual similarity between images. CLIP image score (CLIP-I) [24] measures the semantic similarity between the original and edited images, ensuring the edits align with human perception and maintain contextual relevance. Emotion accuracy (Emo-A) [37] assesses how well the edited image matches the targeted emotion, utilizing a pre-trained emotion classifier. AIM is a more challenging task than emotion generation because it requires both preserving structure and evoking emotions. Therefore, we also introduce a new metric to evaluate the increase in predicted scores for the desired emotion type, called the Emotion Incremental Score (Emo-S). For more details, please refer to the supplementary material.
Figure 5: Ablation study on methodology. Content and color modification, along with ranking technique, are demonstrated to be vital.
@@ 4.2 Comparisons
Since EmoEdit is the first attempt at content editing in AIM, we compare our method with relevant state-of-the-art techniques. These include global editing methods: SDEdit [18], PnP [30], local editing methods: InsDiff [9], P2P-zero [22], ControlNet [42], BlipDiff [15] and style-based AIM methods: CLVA [8], AIF [34].
Qualitative Comparisons
We present the qualitative results in Fig. 8. EmoEdit excels in both preserving structures and evoking emotions compared to other methods. Most compared methods lack sufficient emotion knowledge, resulting in image distortion and severe artifacts. Methods like SDEdit and InsDiff possess a certain level of emotional understanding, attempting to show people with amusement or excitement. However, they struggle to choose contextually appropriate content for the given image and are limited to a single emotion factor. Both CLVA and AIF are emotion style transfer methods that transform realistic images into various artistic styles reflecting different emotions, but they are less flexible in altering the emotional content of the images.
Figure 6: Ablation study on image guidance scale. EmoEdit can progressively edit an input image to different emotion polarities, both positive and negative.
Quantitative Comparisons
In Table 1, EmoEdit outperforms other methods across various metrics. Due to a lack of emotion knowledge, most methods perform poorly on emotion-level metrics, with only 38.07% versus 64.16% in Emo-A, and 0.220 versus 0.363 in Emo-S. One of the biggest challenges in AIM is balancing emotion fidelity with structure integrity. However, aside from emotion-level metrics, EmoEdit also achieves the best results in pixel-level metrics like PSNR and SSIM. For CLIP-I, EmoEdit ranks second, slightly behind PnP, likely because content changes can decrease semantic similarity. Among all compared methods, global editing techniques SDEdit and PnP perform relatively well, possibly because emotion effects are more global rather than local. CLVA and AIF aim to evoke emotions through various artistic styles, effectively preserving image structure but are limited to realistic emotion expression.
Figure 7: EmoEdit can modify a given image across eight different emotional directions, with various factors represented within each specific emotion category.
User Study
We conducted a user study to assess whether humans prefer our method. We invited 54 participants of various ages, with each session lasting about 15 minutes. The study included 40 sets of images, each featuring an original image alongside four edited versions from different methods: SDEdit, P2P-zero, BlipDiffusion, and EmoEdit. Participants were shown a set of images and asked three questions: 1) Which image best preserves the structure? 2) Which image most strongly evokes the targeted emotion? 3) Which image achieves the best balance between structure and emotion? Participants could choose one out of the four options and we calculate the vote percentage for each question. Results in Table 2 show that EmoEdit is the most preferred choice in all questions. Despite the challenge of maintaining structure while conveying emotions, EmoEdit received the highest votes for both aspects. In terms of balance, EmoEdit shows a clear advantage with 87.15% support, confirming that our method aligns with human perceptions.
@@
SDEdit [18] 10.56$\pm$8.0519.17$\pm$8.875.28$\pm$12.07%
P2P-zero [22] 2.64$\pm$2.438.52$\pm$7.661.58$\pm$3.01%
BlipDiff [15] 18.06$\pm$19.6810.46$\pm$6.846.00$\pm$12.02%
EmoEdit 68.75$\pm$23.8661.85$\pm$16.8487.15$\pm$19.89%
@@ 4.3 Ablation study
In Fig. 5, we evaluate the effectiveness of several key network designs in EmoEdit, including content instruction, color instruction and ranking technique. It can be observed that both content and color are crucial visual elements to evoke emotions, with content serving as the foundation and color providing further adjustments. Additionally, ranking technique is essential for selecting the best edit by maintaining structure and evoking emotions. For example, in the case of fear, content instruction replaces flowers with creepy doll heads, and color instruction gives the entire image a gloomy atmosphere. While pumpkins heads are scary, creepy dolls evoke a greater sense of fear.
@@
We show EmoEdit’s editing results with variations in image guidance scale in Fig. 6. The middle image represents the input, while the left and right images illustrate different guidance scales for two emotion polarities: positive and negative. Taking the first image as an example, a race car may evoke excitement, while dead branches and dark lighting can create a sense of fear. We observe that as the image guidance scale decreases, emotion intensity increases while structure preservation diminishes. Although evoking emotion and preserving structure are often contradictory, our method effectively balances them, as demonstrated in Table 1 and Table 2. Users can customize the level of manipulation by adjusting guidance scale to suit their needs and preferences.
@@ 4.4 Applications
Multiple emotion directions
Given an input image, EmoEdit can manipulate it to evoke different emotions by selecting appropriate emotion factors. As illustrated in Fig. 7, taking the window as an example, squirrels may evoke amusement, the aurora elicits awe, fire incites anger and ghost evokes fear. Furthermore, each emotional direction is illustrated with various visual elements, showcasing the diversity of emotion factor trees. For example, emotion factors like surfing on the sea, floating balloons, and beautiful fireworks may evoke excitement, while tombstones, ruined architecture, and shabby houses elicit sadness.
Diverse editing results
In Fig. 9, we present diverse editing results of EmoSet. From an editing perspective, the results can be categorized into low-level (light, color, texture), mid-level (object, scene), and high-level (action and facial expression) changes. From an emotional perspective, the results can be summarized as emotion-generation, emotion-variation, and emotion-enhancement editing. For example, in mid-level editing, adding flowers to the doorway can evoke contentment, turning peaceful dog to fierce tiger can induce anger, and replacing the night sky with majestic canyon can evoke awe. In emotion-enhancement editing, adding balloons can increase amusement, while transforming a grassland under a blue sky into a sea view at sunset brings more awe.
5 Conclusions
We have introduced EmoEdit, a method for evoking emotions by modifying the content and color of images. EmoEdit constructs an emotion factor tree comprising various visual elements. For a given input image, we use VLM and IP2P to generate edit candidates and employ a ranking technique to select the best edits. Our method is evaluated both qualitatively and quantitatively, demonstrating a strong balance between structure preservation and emotion faithfulness. A user study confirms the effectiveness of EmoEdit, and various applications highlight its significant potential.
In the real world, there are many emotion factors beyond those included in our constructed emotion factor tree. Our method relies heavily on EmoSet, so the limited data can introduce some bias and constraints. Exploring a wider range of emotion factors would enhance the richness and effectiveness of the editing results. Additionally, the fixed filter and score ranker in EmoEdit’s ranking technique can limit its effectiveness. Since AIM is a highly human-involved task, incorporating human interaction in future work would provide users with greater flexibility, enabling them to tailor the editing process to their personal tastes and preferences.

### v2

@@ 4.1 Dataset and Evaluation
For fair comparisons, we compile an inference set of 405 images distinct from the training set, sourced from user uploads available online [^1] https://unsplash.com/data. Each image is targeted with 8 emotion directions, resulting in a total of 3,240 image pairs. To capture a broader spectrum of real data, our data comprises positive, negative and neutral images, categorized by a pre-trained emotion classifier on EmoSet [47].
@@
Given the multiple objectives of AIM, we assess our method based on three aspects: pixel-level (PSNR, SSIM), semantic-level (LPIPS, CLIP-I) and emotion-level (Emo-A, Emo-S). PSNR measures the reconstruction quality by comparing the pixel-by-pixel similarity between the original and edited images, indicating the level of noise and distortion introduced during the manipulation process. SSIM [40] evaluates the structural similarity of images by comparing low-level structural information, luminance, and contrast. LPIPS [52] leverages deep learning models to assess the perceptual similarity between images. CLIP image similarity (CLIP-I) [28] measures the semantic consistency between the original and edited images, ensuring the edits align with human perception and maintain contextual relevance. Emotion accuracy (Emo-A) [48] assesses how well the edited image matches the targeted emotion, utilizing a pre-trained emotion classifier [47]. AIM is a more challenging task than emotion generation because it requires both preserving structure and evoking emotions. Therefore, we also introduce a new metric to evaluate the increase in predicted scores for the desired emotion type, called the Emotion Incremental Score (Emo-S). For more details, please refer to the supplementary material.
Figure 6: Ablation study on methodology. Emotion adapter, instruction loss and diffusion loss, are demonstrated to be vital.
Figure 7: Ablation study on image guidance scale. EmoEdit can progressively edit an input image to different emotion polarities.
@@ 4.2 Comparisons
Since EmoEdit is the first attempt at content-aware editing in AIM, we compare our method with the relevant state-of-the-art techniques. These include global editing methods: SDEdit [20], PnP [37], local editing methods: InsDiff [8], P2P-zero [26], ControlNet [51], BlipDiff [15] and style-based AIM methods: CLVA [7], AIF [41].
@@
SDEdit [20] 11.71$\pm$8.9110.85$\pm$7.505.07$\pm$6.08%
P2P-zero [26] 3.05$\pm$3.605.06$\pm$5.930.94$\pm$3.05%
BlipDiff [15] 15.12$\pm$16.138.35$\pm$5.894.88$\pm$10.45%
EmoEdit 70.12$\pm$23.4175.73$\pm$16.4489.12$\pm$14.56%
Qualitative Comparisons
We present the qualitative results in Fig. 5. EmoEdit excels in both preserving structures and evoking emotions compared to other methods. Most compared methods lack sufficient emotion knowledge, resulting in image distortion and severe artifacts. Methods like SDEdit and ControlNet possess a certain level of emotional understanding, attempting to show people with awe or anger. However, they struggle to choose contextually appropriate content for the given image. Both CLVA and AIF are emotion style transfer methods that transform realistic images into artistic styles, but the modifications on different emotions are hard to distinguish.
Quantitative Comparisons
In Table 1, EmoEdit outperforms other methods across various metrics. Due to a lack of emotion knowledge, most methods perform poorly on emotion-level metrics, with only 38.21% versus 50.09% in Emo-A, and 0.221 versus 0.335 in Emo-S. One of the biggest challenges in AIM is balancing emotion fidelity with structure integrity. However, aside from emotion-level metrics, EmoEdit also achieves the best results in pixel-level metrics like PSNR and SSIM. For CLIP-I, EmoEdit ranks second, slightly behind PnP, likely because content changes can decrease semantic similarity. Among all compared methods, global editing techniques SDEdit and PnP perform relatively well, possibly because emotion effects are more global rather than local. CLVA and AIF aim to evoke emotions through artistic styles, which effectively preserves image structure but are restricted to elicit emotions in realistic images.
Figure 8: Emotion adapter can be effectively plugged into existing editing models to enrich their emotional knowledge.
Figure 9: Emotion adapter can be extended to stylized image generation, preserving style and evoking emotions with clear semantics.
User Study
We conducted a user study to assess whether humans prefer our method. We invited 41 participants of various ages, with each session lasting about 15 minutes. The study included 40 sets of images, each featuring an original image alongside four edited versions from different methods: SDEdit, P2P-zero, BlipDiffusion, and EmoEdit. Participants were shown a set of images and asked three questions: (1) Which image best preserves the structure? (2) Which image most strongly evokes the targeted emotion? (3) Which image achieves the best balance between structure and emotion? Participants could choose one out of the four options and we calculate the vote percentage for each question. Results in Table 2 show that EmoEdit is the most preferred choice in all questions. Despite the challenge of maintaining structure while conveying emotions, EmoEdit received the highest votes for both aspects, i.e., 70.12% and 75.73%. In terms of balancing both aspects, EmoEdit shows a clear advantage with 89.12% support, confirming alignment with human perception.
@@ 4.3 Ablation Study
In Fig. 6, we evaluate the effectiveness of several key designs in EmoEdit, including Emotion adapter, instruction loss and diffusion loss. Without the Emotion adapter (w/o EmoAda), images remain nearly identical, highlighting its significance. Both diffusion loss and instruction loss are crucial for evoking emotions. While diffusion loss well preserves the original structure, instruction loss enhances semantic clarity. For instance, in the case of contentment, using diffusion loss adds water to the grass without clear semantics, while using instruction loss adds cakes to the grass, without structural preservation. EmoEdit adds lounge chairs, showcasing structure integrity, semantic clarity and contextual fitting.
@@
We show EmoEdit’s editing results with variations in image guidance scale in Fig. 7. The middle image represents the input, while the left and right images illustrate different guidance scales for two emotion polarities: positive and negative. We observe that as the image guidance scale decreases, emotion intensity increases while structure preservation diminishes. Although evoking emotion and preserving structure are often contradictory, our method effectively balances them, as demonstrated in Table 1 and Table 2. Users can customize the level of manipulation by adjusting guidance scale to suit their preferences.
@@ 4.4 Applications
Once EmoEdit is trained, Emotion adapter can be directly plugged into various diffusion-based models to enhance their emotion awareness with diverse semantics, covering both image editing task and image generation task.
Emotion-enhanced Editing Models
As most existing editing models in Fig. 5 lack emotion knowledge, we experiment to investigate whether Emotion adapter can augment their emotion capabilities in a plug-and-play manner. In Fig. 8, when Emotion adapter is plugged to the original method (w EmoAda), these models are capable of generating images with emotion fidelity and contextual fit. For example, ControlNet transforms the input image into black-and-white, while additional elements, such as thumb stones, are introduced after attaching the Emotion adapter.
Emotion-aware Stylized Image Generation
Apart from editing task, Emotion adapter can also be extended to stylized image generation task in Fig. 9. Once trained, Emotion adapter functions as an emotional interpreter, encoding each emotion polarity (i.e., positive, negative) into distinct semantic representations. We combine the Emotion adapter with Composable Diffusion [18], using 5 art mediums and 5 representative artists as style prompts. Results show that each generated images preserve art styles while evoke specified emotions with clear semantics. Taking Monet as an example, the sunset evokes awe (positive), while the graveyard makes people feel sadness (negative).
5 Conclusion
Discussion
We present EmoEdit, a method designed to evoke emotions by modifying the content of user-provided images. EmoEditSet is first constructed with 40,120 image pairs, serving as a data foundation for AIM. Emotion adapter is proposed and trained using both diffusion loss and the designed instruction loss, which can be directly plugged into diffusion-based generative models to enhance their emotion-awareness. Our method is evaluated both qualitatively and quantitatively, demonstrating a strong balance between structure preservation and emotion faithfulness. User study and ablation study validate the effectiveness of EmoEdit, while applications further highlight the portability of Emotion adapter and the rich emotion knowledge embedded within it.
Limitations
In the real world, there are numerous visual emotion factors beyond those in the constructed emotion factor trees. Besides, due to the inherent complexity of emotions, more than eight emotion categories exist. Our method is heavily dependent on EmoSet, suggesting that the limited data may introduce potential biases and constraints. Expanding the range of emotion factors and categories would improve the quality and diversity of the editing results. Since AIM is a highly human-centered task, future work should place greater emphasis on human interaction. Evaluation metrics, such as emotion accuracy, are largely dependent on the training data and are prone to bias, requiring more human feedback. Additionally, increased flexibility would allow users to tailor the editing process to suit their personal tastes and preferences.